Fermi National Laboratory
 
Operation Summary for November 14, 2011 - November 16, 2011
 

Monday November 14

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the MiniBooNE, NuMI, and the Meson FTBFT-978 experiments taking beam.

During the day shift, Operations turned off the MiniBooNE beam at 9:51 AM to access the MI-12 service building. A MI person began a Booster-MI energy match study at 10:06 AM. At 10:18 AM, the P1 transfer line suffered a vacuum burst when operators reset a vacuum pump; it recovered. Operations turned off the NuMI beam for 15 minutes at 10:24 AM for kicker (I:KPS6N) work. Operations resumed sending beam to MiniBooNE at 1:50 PM. Operators searched and secured the Muon Rings (formerly the Antiproton Source) at 3 PM.

During the evening shift, at 4:32 PM, the P1 transfer line beam permit tripped off when a beam valve (I:BV701) closed; air compressors at MI-20 and MI-60S also tripped off at the same time; operators investigated. Operations reestablished MI beam at 5:29 PM. The MI beam valve, BV701, closed and would not open. Operators thought there was an air leak. This valve problem halted beam to T-978. Operators contacted Mechanical Support personnel and made arrangement to repair the valve early Tuesday morning. Operators reset two P1 transfer line ion pumps at 11:49 PM.

Tuesday November 15

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, Operations stopped sending beam to MI and NuMI at 6 AM in preparation for an access at 7 AM. Operations turned off the MiniBooNE beam at 6:42 AM in preparation for an access later in the morning. At 7 AM, Neutron Therapy Facility personnel began preparing to treat one patient later in the morning. A Mechanical Support person reported at 7:40 AM that they were replacing the airline that goes to BV701.

During the day shift, MiniBooNE personnel accessed MI-65 at 8:35 AM. The MI vacuum work (BV701) ended at 8:54 AM. MiniBooNE personnel came out off MI-65 and operations resumed sending beam to MiniBooNE at 9:25 AM. Operations established beam to T-978 at 9:38 AM. A Muon Ring person began a study at 9:57 AM. Operations established beam to NuMI at 12:22 PM. A MI person began an hour and a half study at 1:01 PM.

During the evening shift, the Muon person completed his study at 4:15 PM. At 4:27 PM, a Linac person reported that he turned off a Klystron RF station (KRF6) due to a LCW leak; a Water Group person repaired the leak and Linac beam returned at 5:16 PM. Operations disabled the T-978 beam at 9:06 PM. Operators completed a tune of the 750 KeV line at 9:53 PM. A Booster RF station (BRF19) tripped off at 10:01 PM; it reset. Operators completed a Booster injection tune at 10:29 PM.

Wednesday November 16

The midnight shift began with Operations sending beam to the NuMI and MiniBooNE.

During the midnight shift, at 1:27 AM, operators reported a Linac 55 LCW system leak; operators investigated and discovered a split hose for a sensor. Operators valved off the leak and contacted the Duty Mechanic. This leak did not interrupt the Linac beam. Operations established beam to T-978 around 7:45 AM.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo approved of the plans by Booster, MI, and Muon personnel to conduct studies. MTA will take beam today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/16/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory